OpenStudy (anonymous):

The equation of line WX is y = -2x - 5. Write an equation of a line perpendicular to line WX in slope-intercept form that contains point (-1, -2).

OpenStudy (anonymous):

perpendicular means that m1*m2=-1 m1=slope of the line WX=-2 m2=slope of the line that we want we could solve for m2

OpenStudy (anonymous):

okay, but whats the answer? sorry im running out of time on my online session and I really just have to do this... :/

OpenStudy (anonymous):

its mult. choice, so these were the options y = -2x y = -1/2 x + 3/2 y= 1/2x -3/2 y=2x

OpenStudy (anonymous):

y= 1/2x -3/2